OpenStudy (anonymous):

I WILL MEDAL/FAN PLEASE, HELP ME. Using the two functions listed below, insert numbers in place of the letters a, b, c, and d so that f(x) and g(x) are inverses. f(x)= x+a b g(x)=cx−d Part 2. Show your work to prove that the inverse of f(x) is g(x).

OpenStudy (anonymous):

\[f(x)=\frac{ x+a }{ b }\]

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

@heretohelpalways @ali2x2

OpenStudy (anonymous):

so is f(x)=x+a over b ok so look at the inverse... swap x and y x=y+a over b can you make y the subject...? ok... multiply both sides by b bx=y+a now you should be able to make y the subject subtract a from both sides well its subtract.. y = bx - a now compare it to the inverse you are given y = cx - d what can you say about b, c and a and d..? it requires you to swap x and y and then make y the subject... so all you need to do, from the way I see it... is to find the inverse of f(x) compare it to g(x) and make the conclusion c = b and a = d then pick numbers for a and b using the conclusion you will also have c and d

OpenStudy (anonymous):

whaaaaattt?

OpenStudy (anonymous):

@pooja195 @ganeshie8